### v2

@@ Appendix A Theoretical Understanding from Data Augmentation Perspective
when $\mathcal{A}(\cdot)$ is instantiated as a mixup function (mixup) that interpolates between the original input data $(x,y)$ and teacher data $(x,y^{t})$ with a mixing coefficient $\lambda=\frac{1}{1+\eta}$, i.e. $y^{\prime}=\lambda y+(1-\lambda)y^{t}$.
@@
where $x_{0}=Y$ and $x_{m+1}=\hat{Y}_{t}^{(m)}$ for $m=0,\dots,M$. By Jensen’s inequality (jensen1906fonctions), for a convex function $f$, we have:
@@
When the temperature $\tau$ goes to $\infty$, minimizing KL divergence between two distributions is equivalent to minimizing the MSE error (kim2021comparing). Thus, we have
@@
and the log sum inequality (thomas2006elements) states that
@@
*The forecastability is borrowed from TimeMixer (timemixer). A larger value indicates better predictability.
@@ Appendix B Implementation Details
Datasets Details. We conduct experiments to evaluate the performance and efficiency of TimeDistill on eight widely used benchmarks: Electricity (ECL), the ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Solar, Traffic, and Weather, following the pipeline in (itransformer; timemixer; moderntcn). Detailed descriptions of these datasets are provided in Table 6.
Metric Details. We use Mean Square Error (MSE) and Mean Absolute Error (MAE) as our evaluation metrics, following (itransformer; moderntcn; timemixer; micn; timesnet; autoformer; informer; dlinear):
@@
Experiment Details. All experiments are implemented in PyTorch (paszke2019pytorch) and conducted on a cluster equipped with NVIDIA A100, V100, and K80 GPUs. The teacher models are trained using their default configurations as reported in their respective papers. For the student MLP model, we employ a combination of a decomposition scheme (dlinear; autoformer; fedformer) and a two-layer MLP with hidden dimension $D$ as the architecture. When using TimeDistill for distillation, the teacher model is frozen, and only the student MLP is trained. The initial learning rate is set to 0.01, and the ADAM optimizer (kingma2014adam) is used with MSE loss for the training of the student model. We apply early stopping with a patience value of 5 epochs. The batch size is set to 32. The temperature for multi-period distillation is set to $\tau=0.5$. The number of scales is set to $M=3$. We use two types of normalization methods, i.e. non-stationary (non-stationary) and revin (revin). We perform a hyperparameter search for $\alpha$ and $\beta$ within the range {0.1, 0.5, 1, 2}. We conduct hyperparameter sensitivity analysis in Appendix H. For consistency, TimeDistill defaults to using ModernTCN as the teacher and fixes the look-back window length to 720. Additional experiments with other teacher models and different look-back window lengths are presented in Section 5.3. Additional detailed model configuration information is presented in Table 7. We also report the standard deviation of TimeDistill performance under five runs with different random seeds in Table 8, which exhibits that the performance of TimeDistill is stable.
@@ Appendix D Further Illustration on Motivation
((itransformer)) ((moderntcn)) ((timemixer)) ((patchtst)) ((micn)) ((fedformer)) ((timesnet)) ((autoformer))
@@ Appendix H Hyperparameter Sensitivity Analysis
iTransformer (itransformer) 0.163 0.258 3.6377 5,276,496
ModernTCN (moderntcn) 0.167 0.262 6.2233 132,075,892
TimeMixer (timemixer) 0.165 0.259 7.7662 5,064,993
PatchTST (patchtst) 0.165 0.266 2.9802 24,949,584
MICN (micn) 0.181 0.293 2.715 60,215,726
Crossformer (crossformer) 0.203 0.299 8.7768 2,264,036
Fedformer (fedformer) 0.274 0.376 32.2642 16,827,399
Autoformer (autoformer) 0.238 0.347 196.5266 14,914,398
@@ Appendix I Directly Match the Patterns in Ground Truth Time Series
TimeDistill may perform suboptimally on time series without strong periodicity or multi-scale patterns. However, such cases are rare, as most real-world time series exhibit these properties (timesnet; autoformer). To verify this, we analyze four datasets using spectral entropy and variance ratio (1H/6H/1D denote different scales). Spectral entropy measures the diversity of periodic components, while variance ratio captures changes across scales. High entropy and decreasing variance ratios indicate strong periodic and multi-scale structure. As shown in the Table 12, real-world datasets show both, unlike synthetic linear series, supporting TimeDistill’s broad applicability.
@@
TimeDistill can be directly adapted to the short-term time series forecasting task. Following TimeMixer (timemixer), we provide some preliminary results in Table 14. We report the results over four PEMS datasets. TimeDistill significantly improves MLP’s performance and surpasses its teacher on the short-term forecasting task.
@@
Our experiment setup follows standard practice (itransformer; non-stationary; sparsetsf) to ensure fairness: each model uses its best-performing normalization based on validation. To reduce confounding factors, we also evaluate all models under the same normalization. As shown in Tables 15 (Non-stationary and REVIN), TimeDistill consistently outperforms prior methods, confirming that its gains come from the distillation method rather than preprocessing differences.
@@
Figure 12. Comparison of model predictions across temporal scales and corresponding spectrograms before and after distillation on the ETTm1 dataset. MSE values for MLP, Teacher (ModernTCN (moderntcn)), and TimeDistill are 2.16, 0.73, and 0.74.
Figure 13. Comparison of model predictions across temporal scales and corresponding spectrograms before and after distillation on the ETTm2 dataset. MSE values for MLP, Teacher (ModernTCN (moderntcn)), and TimeDistill are 3.35, 1.19, and 1.28.
@@
(Ours) ((itransformer)) ((moderntcn)) ((timemixer)) ((patchtst)) ((micn)) ((fedformer)) ((timesnet)) ((autoformer))
@@
Table 23. Ablation study on different datasets (Teacher: iTransformer (itransformer)).
@@
Table 24. Ablation study on different datasets (Teacher: ModernTCN (moderntcn)).

### v3

@@ Appendix A Theoretical Understanding from Data Augmentation Perspective
when $\mathcal{A}(\cdot)$ is instantiated as a mixup function (Zhang, 2017) that interpolates between the original input data $(x,y)$ and teacher data $(x,y^{t})$ with a mixing coefficient $\lambda=\frac{1}{1+\eta}$, i.e. $y^{\prime}=\lambda y+(1-\lambda)y^{t}$.
@@
where $x_{0}=Y$ and $x_{m+1}=\hat{Y}_{t}^{(m)}$ for $m=0,\dots,M$. By Jensen’s inequality (Jensen, 1906), for a convex function $f$, we have:
@@
When the temperature $\tau$ goes to $\infty$, minimizing KL divergence between two distributions is equivalent to minimizing the MSE error (Kim et al., 2021a). Thus, we have
@@
and the log sum inequality (Thomas and Joy, 2006) states that
@@
*The forecastability is borrowed from TimeMixer (Wang et al., 2024a). A larger value indicates better predictability.
@@ Appendix B Implementation Details
Datasets Details. We conduct experiments to evaluate the performance and efficiency of TimeDistill on eight widely used benchmarks: Electricity (ECL), the ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), Solar, Traffic, and Weather, following the pipeline in (Liu et al., 2024; Wang et al., 2024a; Luo and Wang, 2024). Detailed descriptions of these datasets are provided in Table 6.
Metric Details. We use Mean Square Error (MSE) and Mean Absolute Error (MAE) as our evaluation metrics, following (Liu et al., 2024; Luo and Wang, 2024; Wang et al., 2024a, 2023; Wu et al., 2022, 2021; Zhou et al., 2021; Zeng et al., 2023):
@@
Experiment Details. All experiments are implemented in PyTorch (Paszke et al., 2019) and conducted on a cluster equipped with NVIDIA A100, V100, and K80 GPUs. The teacher models are trained using their default configurations as reported in their respective papers. For the student MLP model, we employ a combination of a decomposition scheme (Zeng et al., 2023; Wu et al., 2021; Zhou et al., 2022) and a two-layer MLP with hidden dimension $D$ as the architecture. When using TimeDistill for distillation, the teacher model is frozen, and only the student MLP is trained. The initial learning rate is set to 0.01, and the ADAM optimizer (Kingma, 2014) is used with MSE loss for the training of the student model. We apply early stopping with a patience value of 5 epochs. The batch size is set to 32. The temperature for multi-period distillation is set to $\tau=0.5$. The number of scales is set to $M=3$. We use two types of normalization methods, i.e. non-stationary (Liu et al., 2022) and revin (Kim et al., 2021b). We perform a hyperparameter search for $\alpha$ and $\beta$ within the range {0.1, 0.5, 1, 2}. We conduct hyperparameter sensitivity analysis in Appendix H. For consistency, TimeDistill defaults to using ModernTCN as the teacher and fixes the look-back window length to 720. Additional experiments with other teacher models and different look-back window lengths are presented in Section 5.3. Additional detailed model configuration information is presented in Table 7. We also report the standard deviation of TimeDistill performance under five runs with different random seeds in Table 8, which exhibits that the performance of TimeDistill is stable.
@@ Appendix D Further Illustration on Motivation
((2024)) ((2024)) ((2024a)) ((2023)) ((2023)) ((2022)) ((2022)) ((2021))
@@ Appendix H Hyperparameter Sensitivity Analysis
iTransformer (Liu et al., 2024) 0.163 0.258 3.6377 5,276,496
ModernTCN (Luo and Wang, 2024) 0.167 0.262 6.2233 132,075,892
TimeMixer (Wang et al., 2024a) 0.165 0.259 7.7662 5,064,993
PatchTST (Nie et al., 2023) 0.165 0.266 2.9802 24,949,584
MICN (Wang et al., 2023) 0.181 0.293 2.715 60,215,726
Crossformer (Zhang and Yan, 2023) 0.203 0.299 8.7768 2,264,036
Fedformer (Zhou et al., 2022) 0.274 0.376 32.2642 16,827,399
Autoformer (Wu et al., 2021) 0.238 0.347 196.5266 14,914,398
@@ Appendix I Directly Match the Patterns in Ground Truth Time Series
TimeDistill may perform suboptimally on time series without strong periodicity or multi-scale patterns. However, such cases are rare, as most real-world time series exhibit these properties (2022; 2021). To verify this, we analyze four datasets using spectral entropy and variance ratio (1H/6H/1D denote different scales). Spectral entropy measures the diversity of periodic components, while variance ratio captures changes across scales. High entropy and decreasing variance ratios indicate strong periodic and multi-scale structure. As shown in the Table 12, real-world datasets show both, unlike synthetic linear series, supporting TimeDistill’s broad applicability.
@@
TimeDistill can be directly adapted to the short-term time series forecasting task. Following TimeMixer (Wang et al., 2024a), we provide some preliminary results in Table 14. We report the results over four PEMS datasets. TimeDistill significantly improves MLP’s performance and surpasses its teacher on the short-term forecasting task.
@@
Our experiment setup follows standard practice (Liu et al., 2024, 2022; Lin et al., 2024) to ensure fairness: each model uses its best-performing normalization based on validation. To reduce confounding factors, we also evaluate all models under the same normalization. As shown in Tables 15 (Non-stationary and REVIN), TimeDistill consistently outperforms prior methods, confirming that its gains come from the distillation method rather than preprocessing differences.
@@
Figure 12. Comparison of model predictions across temporal scales and corresponding spectrograms before and after distillation on the ETTm1 dataset. MSE values for MLP, Teacher (ModernTCN (Luo and Wang, 2024)), and TimeDistill are 2.16, 0.73, and 0.74.
Figure 13. Comparison of model predictions across temporal scales and corresponding spectrograms before and after distillation on the ETTm2 dataset. MSE values for MLP, Teacher (ModernTCN (Luo and Wang, 2024)), and TimeDistill are 3.35, 1.19, and 1.28.
@@
(Ours) ((2024)) ((2024)) ((2024a)) ((2023)) ((2023)) ((2022)) ((2022)) ((2021))
@@
Table 23. Ablation study on different datasets (Teacher: iTransformer (Liu et al., 2024)).
@@
Table 24. Ablation study on different datasets (Teacher: ModernTCN (Luo and Wang, 2024)).